A biological sample in a sealed vessel contains a small volume of nitrogen gas at 101.3 kPa at 293.15K. The sample is then stored at -15°C. What is the final pressure of the nitrogen gas?
1
Expert's answer
2020-07-29T06:35:09-0400

Volume pure gass = total volume x partial pressure ratio

P water = 2.34 kPa

P Total = 90.0 kPa

V total = 227ml

P Nitrogen = (90.0 - 2.34) kPa = 87.66kPa

V Nitrogen = Vtotal x partial pressure ratio                                               

227ml x 87.66 / 90 = 221.1 mL    

The volume of pure hydrogen in this sample is 221.1 mL